OpenStudy (anonymous):

Write the quadratic equation in factored form. Be sure to write the entire equation. x^2 + x - 12 = 0

Parth (parthkohli):

Would you like a hint?

OpenStudy (anonymous):

x^2 + x - 12 = 0 x^2 + x+1/4-1/4+12=0 (x+1/2)^2=49/4 x+1/2=7/2 x=7/2-1/2 x=3

OpenStudy (anonymous):